# Bloom: Designing for LLM-Augmented Behavior Change Interactions

### v1

@@ 2.3. Large Language Models and Behavioral Health
Given LLMs offer marked improvements over prior rule-based chatbots in conversational flexibility, recent work has explored their applications to health coaching. Bloom’s LLM health coach is based on the GPTCoach system (Jörke et al., 2025), a GPT-4-based chatbot that implements the onboarding conversation of a physical activity coaching program and uses conversational strategies from MI. Participants reported highly positive and personalized experiences interacting with GPTCoach, and the authors found that it was effective at using strategies aligned with MI principles. Other LLM-based chatbots have explored domains beyond physical activity. Wang et al. (2025) introduced HealthGuru, a sleep health chatbot that uses behavioral guidelines and sensor data to provide recommendations. Xu et al. (2025) developed a chatbot for New Year’s resolutions that personalized suggestions based on chatbot interaction, user profile, and contextual cues (e.g., location, time, goals, etc.). In concurrent work, Heydari et al. (2025) present a multi-agent system consisting of a data science agent, health domain expert agent, and a health coaching agent. Bloom extends these approaches by integrating chat with a suite of UI-based behavior change interactions.
@@
A parallel line of research has focused on improving LLMs’ ability to engage in MI. Studies have applied MI principles in health coaching contexts (Meyer and Elsweiler, 2025; Jörke et al., 2025) and beyond, including mental health (Shah et al., 2022; Chiu et al., 2024), substance use counseling (Kumar et al., 2024; Steenstra et al., 2024), and wellness-related reflection (Rose et al., 2022; Bașar et al., 2025). More recent work has enhanced MI adherence through strategy prediction (Sun et al., 2024) or few-shot demonstrations (Xie et al., 2024). Bloom’s LLM coach conditions each response on a predicted MI strategy.
@@
Beyond chatbots, most studies on LLMs for behavioral health have focused on improving message personalization or sensor data analysis. For example, to improve adherence to the Capacity, Opportunity, Motivation-Behavior (COM-B) framework, Vardhan et al. (2025) explored LLM prompting techniques, whereas Mantena et al. (2025) used a fine-tuning approach to align with the TTM. LLMs have also been applied to downstream health tasks, including generating stage-appropriate behavior change advice (Bak and Chin, 2024), modeling trust and self-referencing in coaching (Meywirth et al., 2025), performing health tasks on raw self-tracking data (Englhardt et al., 2024; Liu et al., 2023; Choube et al., 2025; Li et al., 2025), and extracting insights from personal health data (Subramanian et al., 2024; Imran et al., 2024; Feli et al., 2025; Zhang et al., 2025; Zheng et al., 2025; Khasentino et al., 2025; Fang et al., 2024; Merrill et al., 2024). While these systems show improvement in adhering to theoretical frameworks and personalizing content with quantitative data, they are not designed to proactively elicit and incorporate qualitative information about users’ goals, values, or barriers into a longitudinal coaching process. Bloom explicitly integrates multi-session LLM coaching with behavior change interventions, leveraging the qualitative context elicited from coaching for greater personalization.
@@ 3. Bloom: System Overview
Goal‑setting & Planning 1.1 Goal setting (behavior); 1.2 Problem solving; 1.3 Goal setting (outcome); 1.4 Action planning; 1.5 Review behavior goal(s); 1.6 Discrepancy between current behavior and goal; 1.9 Commitment; 8.7 Graded tasks
Activity Tracking 2.3 Self‑monitoring of behavior; 2.2 Feedback on behavior; 10.4 Social reward
Data Visualizations 1.6 Discrepancy between current behavior and goal; 2.2 Feedback on behavior; 2.3 Self‑monitoring of behavior; 2.4 Self‑monitoring of outcome(s); 2.7 Feedback on outcome(s) of behavior;
Ambient Display 2.2 Feedback on behavior; 7.1 Prompts/cues; 10.3 Non‑specific reward; 10.6 Non-specific incentive; 14.5 Rewarding completion
Push Notifications 1.2 Problem solving; 2.2 Feedback on behavior; 2.3 Self‑monitoring of behavior; 7.1 Prompts/cues; 10.4 Social reward
LLM Chat 1.2 Problem solving; 1.4 Action planning; 1.9 Commitment; 3.1 Social support (unspecified); 4.1 Instruction on how to perform the behavior; 5.1 Information about health consequences; 15.1 Verbal persuasion about capability; 15.3 Focus on past success; 15.4 Self‑talk;
Table 1. Behavior change interactions and the corresponding behavior change techniques (BCTs) they implement, following Michie et al.’s (Michie et al., 2013) taxonomy. While interactions are described at the level of interface components, BCTs capture more granular components underlying effective interventions.
Bloom is an iOS application designed for multi‑week PA support that integrates an LLM coaching agent (Beebo) with UI-based behavior change interactions. In this section, we describe the Bloom system, including the Active Choices program and Bloom’s behavior change interactions, system architecture, LLM coaching agent, and safety filters. Bloom’s user interface is shown in Figures 1 and 2 and our system architecture is shown in Figure 3. Our code is publicly available at https://github.com/StanfordHCI/Bloom.
3.1. Active Choices Program
Bloom implements the Stanford Active Choices Program (King et al., 2002), an evidence-based, scientifically validated counseling program for PA promotion developed by researchers at the Stanford University School of Medicine (Wilcox et al., 2006, 2008; King et al., 2007, 2014; Castro et al., 2011). Active Choices is grounded in the Transtheoretical Model (Prochaska and Velicer, 1997) and Social Cognitive Theory (Bandura, 1999), and many Active Choices facilitators are trained in MI (Miller and Rollnick, 2023). Bloom’s LLM coaching agent directly draws on the program’s structure for onboarding conversations, exercise planning, and check-in conversations. To adapt Active Choices to a digital format, Bloom adds several additional behavior change interactions, described in the following section.
@@ 3.2. Behavior Change Interactions
Bloom integrates the LLM coaching agent with six evidence-based behavior change interactions, leveraging the qualitative context elicited from the agent for richer personalization. In Table 1, we list the behavior change techniques (BCTs), drawn from Michie et al.’s taxonomy (Michie et al., 2013), implemented by each interaction. Whereas interactions are described at the level of the user interface, BCTs capture the underlying “active components” of behavior change interventions and help ground our interactions in techniques known to be effective from behavior change theory.
3.2.1. Goal-Setting & Planning
Beebo, Bloom’s LLM agent, guides the user through weekly, collaborative goal-setting. A weekly goal is an exercise plan that specifies FITT (frequency, intensity, time, type) parameters for each activity (e.g., a moderate-intensity, 20-min walk at 8AM on Mon-Wed-Fri). Beebo first inquires about motivations, long-term goals, past experiences, barriers, resources, and preferences, and subsequently uses this context to propose and confirm commitment to a personalized activity plan. Plans appear in the Plan tab (Figure 2A) and can be edited via the UI or chat. During weekly check-ins, the user’s progress is reviewed. If the user did not meet their goal, barriers are discussed and the plan is revised if necessary. If goals are met and the user remains below recommended activity guidelines, Beebo incrementally suggests longer durations or adding new activity types.
@@ 3.2.2. Activity Tracking
Bloom automatically records activities logged via Apple’s HealthKit API from connected activity trackers. [^1] While our field study required participants to use an Apple Watch, our system supports any wearable that writes to HealthKit. Our system can also support users that do not own a wearable through phone-based tracking and/or manual workout logging. Bloom seeks to link incoming HealthKit workouts to the user’s plan, automatically marking linked activities as complete. Workouts completed outside of a user’s plan are logged as “bonus” activities. Activities that were not recorded in HealthKit can be manually added and/or marked complete.
@@ 3.2.3. Data Visualizations
The Insights tab (Figure 2B) displays visualizations of the user’s data at various granularities, including both metrics directly about behavior (e.g., workouts, step counts) and about outcomes (e.g., resting heart rate). Beebo annotates each chart with natural‑language summaries, describing trends, relating data to the user’s goals, and celebrating progress.
@@ 3.2.4. Ambient Display
Bloom implements a garden-based ambient display (Figure 4) inspired by prior work (Consolvo, 2008; Lin et al., 2006; Murnane et al., 2020). A flower grows in the garden as activities from the weekly plan are completed. Upon achieving their weekly goal, the flower fully blooms and a new flower starts the next week. Bees or butterflies appear above flowers for each completed activity. The ambient display is shown both in the app’s background and on the user’s lockscreen wallpaper (Figure 1E), configured to automatically update via a custom iOS shortcut. Lastly, Beebo generates a congratulatory message (Figure 2C) every time the garden grows. Appendix B provides a full description of the ambient display’s logic.
@@ 3.2.5. Push Notifications
The user receives personalized, LLM-generated notifications from Beebo every morning, every evening, and 15 minutes after each planned activity. Morning notifications remind the user of upcoming activities for the day or celebrate a rest day. If the user has completed the activity, post-activity notifications are congratulatory; otherwise, the notification prompts the user to mark the activity complete or reschedule. Evening notifications celebrate completed activities or ask the user to reflect on their progress.
@@ 3.2.6. LLM Chat
Beebo implements the Active Choices program and uses conversational strategies from MI to offer empathetic and nonjudgmental support, as well as praise for progress. Beebo can brainstorm solutions to common barriers, reschedule or propose alternative activities, and answer PA-related questions.
Figure 2. Additional screens in the Bloom application. (A) The Plan tab shows the user’s current weekly plan, an LLM-generated progress summary, along with missed, upcoming, and past activities. (B) The Insights tab presents visualizations of wearable data, annotated with LLM-generated summaries of trends and progress. (C) When the user’s garden grows, a celebratory modal appears with an LLM-generated message linking progress in the ambient display to recent achievements. (D) During at-will chat, the user can request edits to their plan in natural language, upon which the agent calls plan edit tools. The generated plan is shown as an inline chat widget.

### v2

@@ 2.3. LLMs and Behavioral Health
Given that LLMs offer marked improvements over prior rule-based chatbots in conversational flexibility, recent work has explored their applications to health coaching. Bloom’s LLM health coach is based on the GPTCoach system (Jörke et al., 2025), a GPT-4-based chatbot that implements the onboarding conversation of a PA coaching program and uses conversational strategies from motivational interviewing. Participants reported highly positive and personalized experiences interacting with GPTCoach, and the authors found that it was effective at using strategies aligned with motivational interviewing principles. Other LLM-based chatbots have explored domains beyond PA. Wang et al. (2025) introduced HealthGuru, a sleep health chatbot that uses behavioral guidelines and sensor data to provide recommendations. Xu et al. (2025) developed a chatbot for New Year’s resolutions that personalized suggestions based on chatbot interactions, user profiles, and contextual cues (e.g., location, time, goals, etc.). In concurrent work, Heydari et al. (2025) present a multi-agent system consisting of a data science agent, health domain expert agent, and a health coaching agent. Bloom extends these approaches by integrating chat with a suite of UI-based behavior change interactions.
@@
A parallel line of research has focused on improving LLMs’ ability to engage in motivational interviewing. Studies have applied motivational interviewing principles in health coaching contexts (Meyer and Elsweiler, 2025; Jörke et al., 2025) and beyond, including mental health (Shah et al., 2022; Chiu et al., 2024), substance use counseling (Kumar et al., 2024; Steenstra et al., 2024), and wellness-related reflection (Rose et al., 2022; bașar2025well). More recent work has enhanced motivational interviewing adherence through strategy prediction (Sun et al., 2025) or few-shot demonstrations (Xie et al., 2024). Bloom’s LLM coach conditions each response on a predicted motivational interviewing strategy.
LLMs have also been used to help users make sense of and reflect on their personal health data. Strömel et al. (2024) showed that LLM-generated narratives can foster deeper reflection, more focused attention, and a greater sense of reward compared to conventional charts. Loerakker et al. (2025) study interactions with the commercial WHOOP Coach and identify a “give and take” dynamic in which effective sensemaking requires users to provide sufficient context (give) such that the agent can provide meaningful insight (take). Participants reported mixed impressions, citing generic, impersonal responses and an overly performance-oriented focus as the most common concerns. In contrast, Bloom focuses not only on reflection but on supporting behavior change, using dialogue to elicit context for richer personalization, and employing non-judgmental communication to avoid the pitfalls of performance-oriented support.
@@
Beyond chatbots, most studies on LLMs for behavioral health have focused on improving message personalization or sensor data analysis. For example, to improve adherence to the Capacity, Opportunity, Motivation-Behavior (COM-B) framework, Vardhan et al. (2025) explored LLM prompting techniques, whereas Mantena et al. (2025) used a finetuning approach to align with the TTM. LLMs have also been applied to downstream health tasks, including generating stage-appropriate behavior change advice (Bak and Chin, 2024), modeling trust and self-referencing in coaching (Meywirth et al., 2025), performing health tasks on raw self-tracking data (Englhardt et al., 2024; Liu et al., 2023; Choube et al., 2025; Li et al., 2025), and extracting insights from personal health data (Subramanian et al., 2024; Asif Imran Shouborno et al., 2025; Feli et al., 2025; Zhang et al., 2025; Zheng et al., 2025; Khasentino et al., 2025; Fang et al., 2024; Merrill et al., 2026). While these systems show improvements in adhering to theoretical frameworks and personalizing content with quantitative data, they are not designed to proactively elicit and incorporate qualitative information about users’ goals, values, or barriers into a longitudinal coaching process. Bloom explicitly integrates multi-session LLM coaching with behavior change interventions, leveraging the qualitative context elicited from coaching for greater personalization.
@@ 3. Bloom: System Overview
Figure 2. Additional screens in the Bloom application. (A) The Plan tab shows the user’s current weekly plan, an LLM-generated progress summary, along with missed, upcoming, and past activities. (B) The Insights tab presents visualizations of wearable data, annotated with LLM-generated summaries of trends and progress. (C) When the user’s garden grows, a celebratory modal appears with an LLM-generated message linking progress in the ambient display to recent achievements. (D) During at-will chat, the user can request edits to their plan in natural language, upon which the agent calls plan edit tools, shown as inline widgets.
Additional screens in the Bloom application. A (left): The Plan tab with a weekly plan, showing activities scheduled for the week in a calendar-like widget, an LLM-generated summary congratulating the user, and a list of upcoming and past activities. B (center left): The Insights tab, displaying a weekly summary congratulating the user for their step count and workouts this week, and a bar chart of steps per day. C (center right): A Garden Growth modal, where flowers, butterflies, and bees appear in a garden to celebrate recent workouts, accompanied by a personalized motivational message. D (right): An At-will chat where the user asks Beebo to move a scheduled bike ride, and the agent updates the plan in natural language, responding with an inline plan widget.
Bloom is an iOS application designed for multi-week PA support that integrates an LLM coaching agent (Beebo) with established behavior change interactions. In this section, we describe the Bloom system, including the Active Choices program, Bloom’s behavior change interactions, system architecture, LLM coaching agent, and safety filters. Bloom’s user interface is shown in Figures 1 & 2 and our system architecture is shown in Figure 3. Our code is publicly available at https://github.com/StanfordHCI/Bloom.
3.1. Active Choices Program
Bloom implements the Stanford Active Choices Program (King et al., 2002), an evidence-based, scientifically validated counseling program for PA promotion, typically delivered via phone, developed by researchers at the Stanford University School of Medicine (Wilcox et al., 2006, 2008; King et al., 2007, 2014; Castro et al., 2011). Active Choices is grounded in the Transtheoretical Model (Prochaska and Velicer, 1997) and Social Cognitive Theory (Bandura, 1999). While Active Choices is not a motivational interviewing program, it does advocate for a facilitative communication style and many facilitators are trained in motivational interviewing (Miller and Rollnick, 2023).
Active Choices facilitators follow a structured sequence of conversations with each client, beginning with an onboarding conversation and continuing with regular check-ins. The program provides guidance on communication style and topics to discuss, such as tailoring support to the client’s stage of change, collaboratively setting and reviewing weekly goals, and problem-solving common barriers. It also provides safety considerations drawn from the American College of Sports Medicine (of Sports Medicine, 2025).
Bloom’s LLM coaching agent directly encodes the program’s structure and content for onboarding conversations, PA planning, and check-in conversations in its prompts (described in further detail below). To adapt Active Choices to a digital format, Bloom adds UI elements for planning, activity tracking, and data visualizations. Bloom also introduces interactions not present in the original program, including a garden-based ambient display and push notifications.
@@ 3.1. Active Choices Program
Goal Setting & Planning 1.1 Goal setting (behavior); 1.2 Problem solving; 1.3 Goal setting (outcome); 1.4 Action planning; 1.5 Review behavior goal(s); 1.6 Discrepancy between current behavior and goal; 1.9 Commitment; 8.7 Graded tasks
Activity Tracking 2.3 Self-monitoring of behavior; 2.2 Feedback on behavior; 10.4 Social reward
Data Visualizations 1.6 Discrepancy between current behavior and goal; 2.2 Feedback on behavior; 2.3 Self-monitoring of behavior; 2.4 Self-monitoring of outcome(s); 2.7 Feedback on outcome(s) of behavior;
Ambient Display 2.2 Feedback on behavior; 7.1 Prompts/cues; 10.3 Non-specific reward; 10.6 Non-specific incentive; 14.5 Rewarding completion
Push Notifications 1.2 Problem solving; 2.2 Feedback on behavior; 2.3 Self-monitoring of behavior; 7.1 Prompts/cues; 10.4 Social reward
LLM Chat 1.2 Problem solving; 1.4 Action planning; 1.9 Commitment; 3.1 Social support (unspecified); 4.1 Instruction on how to perform the behavior; 5.1 Information about health consequences; 15.1 Verbal persuasion about capability; 15.3 Focus on past success; 15.4 Self-talk;
Table 1. Behavior change interactions and the corresponding behavior change techniques (BCTs) they implement, following Michie et al.’s (Michie et al., 2013) taxonomy. BCTs capture the psychologically active elements of an intervention and proposed mechanisms of change.
@@ 3.2. Behavior Change Interactions
Bloom integrates the LLM coaching agent with six evidence-based behavior change interactions. Following Slovak & Munson’s framework for psychosocial intervention design in HCI (Slovak and Munson, 2024), we distinguish between the capabilities that LLMs introduce and the intervention components these capabilities enable. We argue that LLMs provide two new capabilities: (1) conducting motivational interviewing conversations with higher fidelity and (2) personalizing intervention content using qualitative context expressed in natural language. We design new intervention components by leveraging these LLM capabilities toward serving a particular psychological function. Some components are drawn directly from Active Choices (e.g., collaborative goal setting), while others augment established components with LLM-driven personalization (e.g., LLM-generated push notifications). This ensemble of components constitutes an intervention system: Bloom.
Below, we list the behavior change interactions implemented by Bloom, which instantiate intervention components using existing UI design patterns and elements. In Table 1, we list the behavior change techniques (BCTs) implemented by each interaction, drawn from Michie et al.’s taxonomy (Michie et al., 2013). BCTs help isolate the psychologically active elements of an intervention and establish the proposed mechanism of change.
3.2.1. Goal Setting & Planning
Beebo, Bloom’s LLM agent, operationalizes the Active Choices program for weekly, collaborative goal setting and planning. By helping users define a meaningful goal, translate it into a concrete weekly plan, and anticipate obstacles, this interaction is intended to increase goal commitment, bolster self-efficacy, and reduce the cognitive load of deciding when and how to act.
In the onboarding conversation, Beebo first introduces the program, confirms the user’s interest and commitment, and proceeds to ask about their high-level motivations, long-term goals, past experiences, barriers, resources, and preferences. Beebo then uses this background information to collaboratively propose, request feedback on, and confirm commitment to a personalized PA plan. The plan is an actionable operationalization of the user’s broader motivations, and defines a short-term goal of completing all planned activities in the coming week. A weekly plan specifies FITT (frequency, intensity, time, type) parameters for each activity (e.g., a moderate-intensity, 20-min walk at 8AM on Mon-Wed-Fri). Plans appear in the Plan tab (Figure 2A) and can be edited via the UI or chat. Consistent with Active Choices, Bloom does not prescribe generic numerical goals such as exercise minutes or steps per day. To ensure plans remain achievable, Beebo is prompted to suggest adjustments if the plan appears overly ambitious. After confirming the plan, Beebo discusses anticipated barriers and, when appropriate, offers advice on overcoming them, and concludes by scheduling the next weekly check-in.
During subsequent weekly check-ins, Beebo and the user review progress using the user’s subjective self-assessment, qualitative reflections, wearable data, and plan completion logs. Following Active Choices, Beebo modulates the difficulty of subsequent plans based on their progress and feedback. If the user completed their plan, Beebo offers praise, explores what helped them succeed, and discusses gently increasing the plan if the user’s current PA levels fall below recommended guidelines for aerobic activity. If the user is already meeting the recommended guidelines, subsequent plans can focus on maintenance or (if appropriate) add strength or flexibility exercises. If the user did not complete their plan, Beebo revisits their motivation and barriers, and proposes adjustments (e.g., simplifying activities, changing timing, or maintaining current levels).
@@ 3.2.2. Activity Tracking
Bloom automatically records activities logged via Apple’s HealthKit API from connected activity trackers. [^1] While our field study required participants to use an Apple Watch, our system supports any wearable that writes to HealthKit. Our system can also support users that do not own a wearable through phone-based tracking and/or manual workout logging. Bloom attempts to link incoming HealthKit workouts to the user’s plan, automatically marking linked activities as complete. Workouts completed outside of a user’s plan are logged as “bonus” activities. Activities that were not recorded in HealthKit can be manually added and/or marked complete. By reducing the effort required to track activity, supporting flexible logging, and recognizing bonus activities, this interaction aims to support self-monitoring and reward progress.
@@ 3.2.3. Data Visualizations
The Insights tab (Figure 2B) displays visualizations of the user’s data at various granularities, including both metrics directly about behavior (e.g., workouts, step counts) and about outcomes (e.g., resting heart rate). Beebo annotates each chart with natural language summaries, and is prompted to describe trends, connect patterns in the data to the user’s goals, and celebrate progress. This interaction aims to help users make sense of their data, offer positive reinforcement, highlight where behavior aligns or diverges from their goals, and support ongoing self-monitoring.
@@ 3.2.4. Ambient Display
Bloom implements a garden-based ambient display (Figure 4) inspired by prior work (Consolvo, 2008; Lin et al., 2006; Murnane et al., 2020). A flower grows in the garden as activities from the weekly plan are completed. Upon achieving their weekly goal, the flower fully blooms and a new flower starts the next week. Bees or butterflies appear above flowers for each completed activity. The ambient display is shown both in the app’s background and on the user’s lockscreen wallpaper (Figure 1E), configured to automatically update via a custom iOS shortcut. Lastly, Beebo generates a congratulatory message (Figure 2C) every time the garden grows. Supplementary Material B provides a full description of the ambient display’s logic. In offering a glanceable, qualitative, and affectively positive representation of a user’s progress, this interaction is intended to offer positive reinforcement and maintain ongoing awareness of activity without requiring deliberate engagement.
Figure 3. System Architecture & Context Management. (A) System Inputs: Bloom draws on wearable data from Apple’s HealthKit API and natural language input from LLM chats. (B) Application Context: Bloom integrates quantitative context, including wearable data and weekly plan progress, with qualitative context in the form of memory summaries from past conversations. (C) LLM Chatbot: Beebo operates in three modes (onboarding, check-in, and at-will chat) and uses dialogue state management and motivational interviewing prompt chains. All responses are passed through safety filters, and the agent can invoke tools to query health data or generate and edit weekly plans. (D) System Outputs: The LLM produces push notifications, natural language plan and data summaries, inline chat widgets in response to tool calls, and celebratory progress messages linked to the ambient display (e.g., garden growth).
System architecture and context management. This figure shows four components. A (left): System Inputs, which include wearable data from Apple HealthKit and chats with Beebo. B (center left): Application context, combining quantitative inputs, including wearable data and weekly plan data, with qualitative context including summary-based memories of past conversation. C (center right): The LLM chatbot, which operates in onboarding, check-in, and at-will modes, using dialogue state management and motivational interviewing prompt chains, and includes safety filters. The agent can also call tools to query health data or create/edit plans. D (right): System outputs, which include push notifications, summaries of plan progress, chat widgets for visualizations and plan edits, and celebratory garden growth messages.
@@ 3.2.5. Push Notifications
The user receives personalized, LLM-generated notifications from Beebo every morning, every evening, and 15 minutes after each planned activity. Morning notifications remind the user of upcoming activities for the day or celebrate a rest day. If the user has completed the activity, post-activity notifications are congratulatory; otherwise, the notification prompts the user to mark the activity complete or reschedule. Evening notifications celebrate completed activities or ask the user to reflect on their progress. By offering timely reminders that either celebrate completed activities or positively reframe missed ones, this interaction supports prospective memory, reduces the effort required for self-monitoring, and helps provide gentle accountability for commitments the user has already made.
@@ 3.2.6. LLM Chat
Beebo implements the Active Choices program and uses conversational strategies from motivational interviewing to offer empathetic and non-judgmental support. Beebo can brainstorm solutions to common barriers, reschedule or propose alternative activities, and answer PA-related questions. Beebo serves multiple psychological functions as the delivery channel for several aforementioned components (e.g., goal setting, planning, problem-solving barriers). Moreover, Beebo’s implementation of motivational interviewing strategies aims to promote the users’ sense of autonomy, motivation, and self-efficacy.